# Taxonomy and Estimation of Multiple Breakpoints in High-Dimensional Factor Models

### v2

@@ 4.1 Assumptions
There exists an $M<\infty$ such that for all values of $N$ and $T$,
(i) $E\left(\max_{\begin{subarray}{c}k-s>(\log T)^{2}\end{subarray}}\frac{1}{k-s}\sum_{t=s+1}^{k}\left\|\frac{1}{\sqrt{NT}}\sum_{h=1}^{T}\sum_{i=1}^{N}f_{h}[e_{ih}e_{it}-E(e_{ih}e_{it})]\right\|^{2}\right)\leq M;\\$
(ii) $\lim\sup\limits_{N\to\infty,k-s>(\log T)^{2}}\left\|\frac{1}{\sqrt{N(k-s)\log\log(N(k-s))}}\sum\limits_{t=s+1}^{k}\sum\limits_{i=1}^{N}\lambda_{i}f_{t}^{\prime}e_{it}\right\|^{2}\leq M;$
(iii) $\lim\sup\limits_{N\to\infty,k-s>(\log T)^{2}}\left\|\frac{1}{N\sqrt{k-s}\sqrt{\log\log N^{2}(k-s)}}\sum\limits_{t=s+1}^{k}\sum\limits_{i=1}^{N}\sum\limits_{h=1}^{N}\lambda_{i}\lambda_{h}^{{}^{\prime}}(e_{it}e_{ht}-E(e_{it}e_{ht}))\right\|\leq M.$
Assumptions 1-5 are standard in the factor model literature. Assumption 1 (i) is similar to Assumption A in Bai (2003) and (ii) imposes restrictions on $B_{j}$ and $B_{j-1}$ to ensure the identification of the breakpoint. Assumption 1 (ii) rules out the case that $B_{j}=-B_{j-1}$ since the objective function (18) can not identify the difference between $\hat{\Sigma}_{j-1}$ and $\hat{\Sigma}_{j}$. Assumption 1 (iii) implies that $\Sigma_{G}$ is positive definite. Assumption 2 is similar to Assumption B of Bai (2003). Assumption 3 allows for weakly correlated error terms in both time and cross-sectional dimensions. Assumption 4 means that the factors and idiosyncratic errors are allowed to be weakly dependent within each regime. Assumption 5 corresponds to Assumption G in Bai (2003). Assumption 6 corresponds to Assumption 7 of Baltagi et al. (2021), which allows the Hájek-Rényi inequality to be applicable to the second moment process of the factors. Assumptions 7 and 8 are similar to Assumptions 7 and 8 of DBH. Assumptions 7 and 8(i) impose further conditions on the weak dependence among the idiosyncratic error, while parts (ii) and (iii) of Assumption 8 are implications of the Law of the Iterated Logarithm (LIL).
@@
when $B_{j+1}$ is singular or $\|B_{j+1}f_{k_{j}^{0}+1}-\mathrm{Proj}(B_{j+1}f_{k_{j}^{0}+1}|B_{j})\|\geq d>0$ when $B_{j}$ is singular, where $\mathrm{Proj}(\mathbb{A}|\mathbb{Z})$ denotes the projection of $\mathbb{A}$ onto the columns of $\mathbb{Z}$, and $d$ is a constant.
Assumption 9 extends Assumption F3 in Bai (2003) and ensures that $\rho_{r}(\hat{\Sigma}(k,s))\asymp\frac{1}{N}$ w.p.a.1 if the interval $[k+1,s]$ belongs to $j$-th regime and $B_{j}$ is singular. Assumption 10 rules out certain degenerate cases where the projection error is zero. Specifically, it prevents $f_{k_{j}^{0}}$ from lying in the null space of $B_{j}$ (or $f_{k_{j}^{0}+1}$ in the null space of $B_{j+1}$) and also ensures that $f_{k_{j}^{0}}$ and $f_{k_{j}^{0}+1}$ are not zero. Either scenario would result in a zero projection error, violating the assumption.
@@ 4.2 Asymptotic properties of the QML estimators
In this subsection, we establish the asymptotic properties of the QML estimators. The analysis of these properties is significantly more challenging in a multi-break setting than in a single-break setting. As discussed in the Introduction, the estimation error of one estimated breakpoint is generally interdependent with that of others. This interdependence is further complicated when both Type 1 and Type 2 breaks are present. Furthermore, since the true endpoints of a regime are unknown, an estimated segment $\hat{k}_{j}+1,\hat{k}_{j+1}]$ may have no overlap with the true interval $k_{j}^{0}+1,k_{j+1}^{0}]$. This contrasts sharply with the single-break setting, where the estimated segments are guaranteed to overlap with the true regimes because $k_{0}^{0}=1$ and $k_{m_{0}+1}^{0}=k_{2}^{0}=T$ are known.
To address these challenges, we introduce a strategy of overfitting the number of breaks to isolate the $j$-th breakpoint of interest. Specifically, assuming $\hat{k}_{j}\neq k_{j}^{0}$, we conceptually insert all true breakpoints, except for the $j$-th one, into the existing set of estimated breakpoints. After inserting these breakpoints, all resulting segments (derived from the estimated and inserted breakpoints) contain no true breakpoint, except for one segment which contains the $j$-th true breakpoint. This enables us to derive an upper bound for $|\hat{k}_{j}-k_{j}^{0}|$.
While this strategy helps isolate the target breakpoint and analyze its asymptotic properties, the overfitting strategy introduces two technical challenges. First, an inserted true breakpoint may be very close to an estimated breakpoint, leading to a segment with fewer than $r$ observations. In such a segment, the pseudo-factors will have an exactly singular covariance matrix. We address this issue by merging such a short segment with an adjacent regime whose length is necessarily proportional to $T$. More details are provided in the text preceding the proof of Theorem 1 in the Appendix. Second, overfitting the breaks inevitably cuts a true regime into sub-regimes. When the pseudo-factors in the $j$-th regime are not of full column rank (i.e., $B_{j}$ is singular), it is crucial and challenging to compare the nearly zero eigenvalues of the sample covariance matrices of pseudo-factors in the true regimes versus those in the sub-regimes created by the overfitting. Proposition 2 provides a detailed rate for the difference between these nearly-zero eigenvalues. This proposition plays a crucial role in establishing the point-consistency of the QML estimator and the selection consistency of our proposed information criterion.
To compare these nearly-zero eigenvalues, we consider the case of overfitting a point $\tilde{k}_{j}$ within the regime $k_{j-1}^{0}+R,\ k_{j}^{0}-R]$, where $R=(\log T)^{2}$. Recall that $\hat{\Sigma}(k,s)$ denotes the estimated sample covariance matrix of $\hat{g}_{t}$ for $k+1\leq t\leq s$. Specifically,
@@
Under Assumptions 1-8 and $\frac{N}{T}\to\kappa$, as $N,\,T\to\infty$ for $0<\kappa<\infty$, when $\mathrm{rank}(B_{j})=r_{j}<r$,
@@
Theorem 1 provides an initial upper bound on the estimation error of $\hat{k}_{j}$ for both types of breaks. Part (i) indicates that the estimation error for Type 1 breaks is bounded by $\sqrt{T}$ with high probability in large samples. Moreover, Part (ii) shows that for a Type 2 break, the upper bound is $T^{\alpha}$, where $\alpha$ is some constant slightly larger than $1/2$. These results collectively imply fraction-consistency, i.e., $\tau_{j}-\tau_{j}^{0}=o_{p}(1)$, where $\tau_{j}=\hat{k}_{j}/T$ is the estimated break fraction. Building upon Theorem 1, the subsequent theorem further refines the convergence rate of $\hat{k}_{j}-k_{j}^{0}$.
@@
Theorem 2 (i) demonstrates that the estimated change points converge to the true breakpoints w.p.a.1 for a Type 1 break. This result is stronger than that of Baltagi et al. (2021), who establish that the distance between estimated and true breakpoints is $O_{p}(1)$ for Type 1 changes. It is important to note that in Type 1 scenarios, where $r_{j,j+1}=\max\{r_{j},r_{j+1}\}$, the model accounts for the emergence and disappearance of factors. Our QML estimator proves to be consistent when the number of factors changes acorss regimes, whereas Ma and Su (2018) exclude this case by assumption. Furthermore, Theorem 2 (ii) suggests that the discrepancy between the QML estimators and the true change points is stochastically bounded for Type 2 changes.
To validate these theoretical findings, we conducted simulations involving factor loadings with rotational changes (see DGP 1.D and DGP 1.E in Section 5). The simulation results reveal that our QML estimators exhibit significantly lower MAEs and RMSEs compared to those reported by Baltagi et al. (2021) and Ma and Su (2018).
@@
We discuss how to analyze the nature of changes around the breakpoints and assess the reliability of the estimated breakpoint. For clarity, let us denote the estimated breakpoints as $\hat{k}_{j-1}$, $\hat{k}_{j}$, and $\hat{k}_{j+1}$, with the aim of determining the type of change around $\hat{k}_{j}$. In practical applications, by utilizing our QML estimators and following Bai and Ng’s (2002) information criterion, one can estimate the number of factors $r_{j}$, $r_{j+1}$, and $r_{j,j+1}$ for the regimes $\hat{k}_{j-1}+1,\hat{k}_{j}]$, $\hat{k}_{j}+1,\hat{k}_{j+1}]$, and the combined regime $\hat{k}_{j-1}+1,\hat{k}_{j+1}]$, respectively. (A) If $r_{j}=r_{j+1}=r_{j,j+1}$, we classify the breakpoint as a rotational change (Type 2). Within this category, if $r_{j,j+1}=r$, the change across the breakpoint is a full-rank rotational change. Conversely, if $r_{j,j+1}<r$, the change remains rotational, but it indicates that the $B_{j}$ matrices are singular. This singularity is not induced by the rotational change itself; rather, it reflects the presence of other singular changes in the overall sample. (B) If $r_{j,j+1}>\min\{r_{j},r_{j+1}\}$, the $j$-th break is a singular change (Type 1). In this category: if $r_{j,j+1}=r_{j}+r_{j+1}$, the columns of the factor loading matrices before and after the breakpoint are linearly independent—a scenario commonly explored in existing literature. If $r_{j,j+1}=r_{j}>r_{j+1}$ (or $r_{j,j+1}=r_{j+1}>r_{j}$), the columns of factor loading matrix after (before) the breakpoint form a subspace of those before (after) the breakpoint. This means some factors disappear (or some new factors emerge) after the breakpoint. For case (A), the difference between the estimated and true breakpoint is stochastically bounded, as shown by Theorem 2 (ii). In contrast, for case (B), the estimated breakpoints are equal to the true breakpoints w.p.a.1, as indicated by Theorem 2 (i). Therefore, one can assess the type of break around the breakpoint and evaluate the reliability of the estimated breakpoint based on values of $r$, $r_{j,j+1}$, $r_{j}$, and $r_{j+1}$.
@@
$IC(m)$ $=$ $\sum\limits_{\ell=1}^{m+1}(k_{\ell}-k_{\ell-1})\log|\hat{\Sigma}(k_{\ell-1},k_{\ell})|+m(1+|\hat{\rho}|)r^{2}\log(\min(N,T)),$ (30)
where $\hat{\rho}$ is the radius of the estimated AR(1) coefficient matrix, obtained by fitting a VAR(1) model to $\hat{g}_{t}$. Let $m_{max}$ be a bounded integer such that $m_{0}<m_{max}$ and $\hat{m}=\arg\min_{1\leq m\leq m_{max}}IC(m)$.

### v3

@@ 4.1 Assumptions
The following conditions hold for some finite constant $M>0$:
(i) For a fixed $s$ and varying $k$ (or vice-versa), such that $k-s>(\log T)^{2}$:
$\limsup_{N,(k-s)\to\infty}\left\|\frac{1}{\sqrt{N(k-s)\log\log(N(k-s))}}\sum_{t=s+1}^{k}\sum_{i=1}^{N}\lambda_{i}f_{t}^{\prime}e_{it}\right\|\leq M;$
And for any constant fraction $\eta\in(0,1]$, let $\mathcal{S}_{\eta}=\{(s,k):k-s>\eta T\}$. Then
$\sup_{(s,k)\in\mathcal{S}_{\eta}}\left\|\frac{1}{\sqrt{N(k-s)}}\sum_{t=s+1}^{k}\sum_{i=1}^{N}\lambda_{i}f_{t}^{\prime}e_{it}\right\|=O_{p}(1);$
(ii) For a fixed $s$ and varying $k$ (or vice-versa), such that $k-s>(\log T)^{2}$:
$\limsup_{N,(k-s)\to\infty}\left\|\frac{1}{N\sqrt{k-s}\sqrt{\log\log(N^{2}(k-s))}}\sum_{t=s+1}^{k}\sum_{i=1}^{N}\sum_{j=1}^{N}\lambda_{i}\lambda_{j}^{\prime}(e_{it}e_{jt}-E[e_{it}e_{jt}])\right\|\leq M;$
And for any constant fraction $\eta\in(0,1]$, let $\mathcal{S}_{\eta}=\{(s,k):k-s>\eta T\}$. Then
$\sup_{(s,k)\in\mathcal{S}_{\eta}}\left\|\frac{1}{N\sqrt{k-s}}\sum_{t=s+1}^{k}\sum_{i=1}^{N}\sum_{j=1}^{N}\lambda_{i}\lambda_{j}^{\prime}(e_{it}e_{jt}-E[e_{it}e_{jt}])\right\|=O_{p}(1).$
Assumptions 15 are standard in the factor model literature. Assumption 1 (i) is similar to Assumption A in Bai (2003) and (ii) imposes restrictions on $B_{j}$ and $B_{j-1}$ to ensure the identification of the breakpoint. Assumption 1 (ii) rules out the case that $B_{j}=-B_{j-1}$ since the objective function (18) cannot identify the difference between $\hat{\Sigma}_{j-1}$ and $\hat{\Sigma}_{j}$. Our analysis focuses on the case of fixed-magnitude breaks to facilitate the development of the statistical properties of near-zero eigenvalues established in Proposition 2. Assumption 1 (iii) implies that $\Sigma_{G}$ is positive definite. Assumption 2 is similar to Assumption B of Bai (2003). Assumption 3 allows for weakly correlated error terms in both time and cross-sectional dimensions. Assumption 4 means that the factors and idiosyncratic errors are allowed to be weakly dependent within each regime. Assumption 5 corresponds to Assumption G in Bai (2003). Assumption 6 corresponds to Assumption 7 of Baltagi et al. (2021), which allows the Hájek–Rényi inequality to be applicable to the second moment process of the factors.
Assumptions 7 and 8 are similar to Assumptions 7 and 8 of DBH. The terms within the squared norm in Assumption 7 are standard in the factor model literature and are typically $O_{p}(1)$ (e.g., Assumptions C(5), F(1), and F(3) in Bai (2003)). Our assumption strengthens these traditional conditions by requiring that the averages of these squared terms remain uniformly $O_{p}(1)$ over sub-intervals with length exceeding $(\log T)^{2}$. We provide sufficient conditions for this uniformity in Section C of the Supplementary Appendix. Assumption 8 concerns the partial sums of the mean-zero processes $Z_{N,t}^{(1)}$ and $Z_{N,t}^{(2)}$ over candidate sub-intervals $(s,k]$, where $Z_{N,t}^{(1)}=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\lambda_{i}f_{t}^{\prime}e_{it}$ and $Z_{N,t}^{(2)}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}\lambda_{i}\lambda_{j}^{\prime}\bigl(e_{it}e_{jt}-E[e_{it}e_{jt}]\bigr)$. For $k-s\geq\eta T$, the required $O_{p}(1)$ bound is a direct consequence of a Functional Central Limit Theorem. For shorter intervals where $k-s\geq(\log T)^{2}$, the relevant upper envelope involves a $\log\log$ factor, consistent with the Law of the Iterated Logarithm. Detailed derivations for a benchmark example illustrating the feasibility of Assumption 8 are provided in Section C of the Supplementary Appendix.
@@
when $B_{j+1}$ is singular or $\|B_{j+1}f_{k_{j}^{0}+1}-\mathrm{Proj}(B_{j+1}f_{k_{j}^{0}+1}|B_{j})\|\geq d>0$ when $B_{j}$ is singular, where $\mathrm{Proj}(\cdot|\mathbb{Z})$ denotes the projection of a vector onto the columns of $\mathbb{Z}$, and $d$ is a constant.
Assumption 9 extends Assumption F3 in Bai (2003) and ensures that $\rho_{r}(\hat{\Sigma}(k,s))\asymp\frac{1}{N}$ w.p.a.1 if the interval $[k+1,s]$ belongs to $j$-th regime and $B_{j}$ is singular. Assumption 10 is critical when $B_{j}$ or $B_{j+1}$ is singular. Suppose $B_{j+1}$ is singular with $r_{j+1}<r$. Under the true partition, the pseudo-factor covariance matrix for $(k_{j}^{0},k_{j+1}^{0}]$ has rank $r_{j+1}$. Shifting the split point from $k_{j}^{0}$ to $k_{j}^{0}-1$ mistakenly incorporates $B_{j}f_{k_{j}^{0}}$ from regime $j$ into the misaligned segment $(k_{j}^{0}-1,k_{j+1}^{0}]$. Assumption 10 requires $B_{j}f_{k_{j}^{0}}\notin\mathrm{Col}(B_{j+1})$, ensuring that the misaligned segment’s rank exceeds $r_{j+1}$. In finite samples, where estimated factors are used to compute $\hat{\Sigma}(k_{\ell-1},k_{\ell})$, this rank inflation translates into an increase in the objective function, facilitating precise breakpoint identification. Conversely, if $B_{j}f_{k_{j}^{0}}\in\mathrm{Col}(B_{j+1})$ (e.g., $f_{k_{j}^{0}}=0$), misspecifying the breakpoint as $k_{j}^{0}-1$ fails to increase the rank of segment $(k_{j}^{0}-1,k_{j+1}^{0}]$, in which case the breakpoint cannot be precisely estimated. The case of $B_{j}$ being singular is similar by symmetry.
@@ 4.2 Asymptotic properties of the QML estimators
In this subsection, we establish the asymptotic properties of the QML estimators. The analysis of these properties is significantly more challenging in a multi-break setting than in a single-break setting. As discussed in the Introduction, the estimation error of any estimated breakpoint is generally interdependent with that of others. This interdependence is further complicated when both Type 1 and Type 2 breaks are present. Furthermore, since the true endpoints of a regime are unknown, an estimated segment $\hat{k}_{j}+1,\hat{k}_{j+1}]$ may have no overlap with the true interval $k_{j}^{0}+1,k_{j+1}^{0}]$. This contrasts sharply with the single-break setting, where the estimated segments are guaranteed to overlap with the true regimes because $k_{0}^{0}=0$ and $k_{m_{0}+1}^{0}=k_{2}^{0}=T$ are known.
To address these challenges, we introduce a strategy of “conceptually” overfitting the number of breaks to isolate the $j$-th breakpoint of interest. Specifically, assuming $\hat{k}_{j}\neq k_{j}^{0}$, we conceptually insert all true breakpoints, except for the $j$-th one, into the existing set of estimated breakpoints. After inserting these breakpoints, all resulting segments (derived from the estimated and inserted breakpoints) contain no true breakpoint, except for one segment that contains the $j$-th true breakpoint. This enables us to derive an upper bound for $|\hat{k}_{j}-k_{j}^{0}|$.
While this strategy helps isolate the target breakpoint and analyze its asymptotic properties, the conceptually overfitting strategy introduces two technical challenges. First, an inserted true breakpoint may be very close to an estimated breakpoint, leading to a segment with fewer than $r$ observations. In such a segment, the pseudo-factors will have an exactly singular covariance matrix. We address this issue by merging such a short segment with an adjacent regime whose length is necessarily proportional to $T$. More details are provided in the text preceding the proof of Theorem 1 in the Appendix. Second, overfitting the breaks inevitably cuts a true regime into sub-regimes. When the pseudo-factors in the $j$-th regime are not of full column rank (i.e., $B_{j}$ is singular), it is crucial and challenging to compare the nearly zero eigenvalues of the sample covariance matrices of pseudo-factors in the true regimes versus those in the sub-regimes created by conceptually overfitting. Proposition 2 provides a detailed rate for the difference between these nearly-zero eigenvalues. This proposition plays a crucial role in establishing the point-consistency of the QML estimator and the selection consistency of our proposed information criterion.
To compare these nearly-zero eigenvalues, we consider the case of conceptually overfitting a point $\tilde{k}_{j}$ within the regime $k_{j-1}^{0}+R,\ k_{j}^{0}-R]$, where $R=(\log T)^{2}$. Recall that $\hat{\Sigma}(k,s)$ denotes the estimated sample covariance matrix of $\hat{g}_{t}$ for $k+1\leq t\leq s$. Specifically,
@@
Under Assumptions 18 and $\frac{N}{T}\to\kappa$, as $N,\,T\to\infty$ for $0<\kappa<\infty$, when $\mathrm{rank}(B_{j})=r_{j}<r$, then
@@
Theorem 1 provides an initial upper bound on the estimation error of $\hat{k}_{j}$ for both types of breaks. Part (i) indicates that the estimation error for Type 1 breaks is bounded by $\sqrt{T}$ with high probability in large samples. Moreover, Part (ii) shows that for a Type 2 break, the upper bound is $T^{\alpha}$, where $\alpha$ is some constant slightly larger than $1/2$. These results collectively imply fraction-consistency, i.e., $\hat{\tau}_{j}-\tau_{j}^{0}=o_{p}(1)$, where $\hat{\tau}_{j}=\hat{k}_{j}/T$ is the estimated break fraction. Building upon Theorem 1, the subsequent theorem further refines the convergence rate of $\hat{k}_{j}-k_{j}^{0}$.
@@
Theorem 2 (i) demonstrates that the estimated change points converge to the true breakpoints w.p.a.1 for a Type 1 break. This result is stronger than that of Baltagi et al. (2021), who established that the distance between estimated and true breakpoints is $O_{p}(1)$ for Type 1 changes. It is important to note that Type 1 breaks incorporate scenarios of emergence and disappearance of factors, which correspond to $r_{j,j+1}=\max\{r_{j},r_{j+1}\}$. Our QML estimator proves to be consistent when the number of factors changes across regimes, whereas Ma and Su (2018) excluded this case by assumption. Furthermore, Theorem 2 (ii) suggests that the discrepancy between the QML estimators and the true change points is stochastically bounded for Type 2 changes.
@@
[…]
Remark 6 (Determining the types of changes)
We discuss how to analyze the nature of changes around the breakpoints and assess the reliability of the estimated breakpoint. To implement the classification rule below in practice, the ranks $(r_{j},r_{j+1},r_{j,j+1})$ can be estimated from the data using the step-by-step procedure described in Remark 7. (A) If $r_{j}=r_{j+1}=r_{j,j+1}$, we classify the breakpoint as a rotational change (Type 2). This includes both the full-rank case $r_{j,j+1}=r$ and the case $r_{j,j+1}<r$ (see Remark 3). (B) If $r_{j,j+1}>\min\{r_{j},r_{j+1}\}$, the $j$-th break is a singular change (Type 1). If $r_{j,j+1}=r_{j}+r_{j+1}$, the loading spaces before and after the breakpoint are linearly independent. If $r_{j,j+1}=r_{j}>r_{j+1}$ (or $r_{j,j+1}=r_{j+1}>r_{j}$), the loading space after (before) the breakpoint is a subspace of that before (after) the breakpoint, implying that some factors disappear (or new factors emerge) after the breakpoint. For case (A), Theorem 2(ii) shows that the estimation error is stochastically bounded. For case (B), Theorem 2(i) shows that the estimated breakpoints equal the true breakpoints w.p.a.1. Therefore, one can determine the type of break and assess the reliability of the estimated breakpoint based on the values of $r$, $r_{j}$, $r_{j+1}$, and $r_{j,j+1}$.
Remark 7 (Estimating $(r_{j},r_{j+1},r_{j,j+1})$ in practice)
The number of pseudo-factors in full sample, r, is consistently estimated using the information criteria (IC) of Bai and Ng (2002). Given $\hat{r}$, we obtain the estimated breakpoints $\{\hat{k}_{1},\dots,\hat{k}_{m_{0}}\}$. To implement the classification rule in Remark 6, we apply the Bai and Ng (2002) IC—with $\hat{r}$ as the upper bound—to three specific subsamples: $\hat{k}_{j-1}+1,\hat{k}_{j}]$, $\hat{k}_{j}+1,\hat{k}_{j+1}]$, and $\hat{k}_{j-1}+1,\hat{k}_{j+1}]$. Let the resulting rank estimates be $\hat{r}_{j}$, $\hat{r}_{j+1}$, and $\hat{r}_{j,j+1}$, respectively. Theorem 2 establishes that $\hat{k}_{j}$ is point-consistent under Type 1 changes and satisfies $\hat{k}_{j}-k_{j}^{0}=O_{p}(1)$ under Type 2 changes. Since the distance between true breakpoints is of order $T$, the $O_{p}(1)$ estimation error in the breakpoints has a negligible effect on rank estimation; thus, Bai and Ng’s IC remains consistent for these subsamples. To ensure finite-sample robustness, we apply the adjustment $\tilde{r}_{j,j+1}=\max\{\hat{r}_{j,j+1},\hat{r}_{j},\hat{r}_{j+1}\}$, which guarantees $\max\{\hat{r}_{j},\hat{r}_{j+1}\}\leq\tilde{r}_{j,j+1}\leq\hat{r}$ by construction. The breakpoint type is then determined via $(\hat{r}_{j},\hat{r}_{j+1},\tilde{r}_{j,j+1})$ according to Remark 6.
@@
$IC(m)$ $=$ $\sum\limits_{\ell=1}^{m+1}(\hat{k}_{\ell}-\hat{k}_{\ell-1})\log|\hat{\Sigma}(\hat{k}_{\ell-1},\hat{k}_{\ell})|+m(1+|\hat{\rho}|)r^{2}\log(\min(N,T)),$ (30)
where $(\hat{k}_{1},\dots,\hat{k}_{m})$ are obtained by solving (18) for each given $m$, and $\hat{\rho}$ is the radius of the estimated AR(1) coefficient matrix, obtained by fitting a VAR(1) model to $\hat{g}_{t}$, although the actual $g_{t}$ is not required to be an AR(1) process.
Note that, for any fixed $m$, the penalty term in Eq. (30) is constant with respect to the breakpoint locations. Hence, the joint search over $m$ and the breakpoint locations is numerically equivalent to the following two-step procedure: first, for each $m$, determine the optimal break locations by minimizing the QML objective function in Eq. (18); second, substitute these minimizers into Eq. (30) and select the value of $m$ that yields the smallest criterion value. In practice, for each candidate $m$, the minimization of $U_{NT}$ is implemented by the dynamic programming algorithm in Section 2.2.
Let $m_{max}$ be a bounded integer such that $m_{0}<m_{max}$ and $\hat{m}=\arg\min_{0\leq m\leq m_{max}}IC(m)$.
@@
Remark 8
Theorems 2 and 3 together imply the following joint consistency result:
[…]
where $S=\{j\mid r_{j,j+1}>\min\{r_{j},r_{j+1}\},\ j=1,\ldots,m_{0}\}$ denotes the set of indices for singular breaks. If we also include rotational changes, then for any given $\epsilon>0$ there exists an $M>0$ such that $P(|\hat{k}_{j}-k_{j}^{0}|<M,\,j=1,2,\cdots,m_{0},\ \hat{m}=m_{0})>1-\epsilon$.